# AdaptiveK Sparse Autoencoders: Dynamic Sparsity Allocation for Interpretable LLM Representations

### v1

@@ 4.3 Pareto Frontier Results
We evaluated the three different metrics vs. sparsity frontier for benchmark SAEs with different sparsity constraints. All SAEs have a dictionary size of 16384. These three metrics are:
@@
  • Reconstruction L2 Loss: $\|x-\hat{x}\|_{2}^{2}$ averages the squared Euclidean distance between each reconstruction and its input, the smaller the better.
@@
  • Fraction of Variance Explained: $\frac{\mathrm{Var}(x-\hat{x})}{\mathrm{Var}(x)}$ measures how much input variability is captured by the reconstruction. The plots show 1 minus this value (unexplained variance, the smaller the better).
@@
  • Cosine Similarity: $\frac{x\cdot\hat{x}}{\|x\|_{2}\|\hat{x}\|_{2}}$ evaluates directional fidelity between original and reconstructed vectors. The plots show 1 minus this value, thus the smaller the better.
Figure 5, 6 and 7 demonstrate that across Pareto Frontier results on different LLMs, AdaptiveK’s reconstruction error remains consistently lower than all other SAE approaches, even as they improve with increased sparsity. For cosine similarity and explained variance, AdaptiveK outperforms at equivalent sparsity levels. While some SAEs match or exceed these metrics at extremely high sparsity (over ten times greater), such cases violate the sparsity-fidelity tradeoff beacuse an infinite-width SAE can theoretically reconstruct perfectly (Karvonen et al., 2025). Thus, our SAE transcends the traditional Pareto Frontier, achieving results unreachable by other SAEs regardless of parameter tuning.
@@
Notably, across all three model scales, AdaptiveK’s reconstructed activations more accurately match the originals in both distance and direction (Figure 5(b), 6(b), 7(b), 5(c), 6(c), 7(c)) while explaining a higher proportion of data variance (Figure 5(a), 6(a), 7(a)). This consistent cross-model performance demonstrates AdaptiveK’s robustness and generalizability.
(a) Sparsity vs. Unexplained
(b) Sparsity vs. L2 Loss
(c) Sparsity vs. Cosine Similarity
Figure 5: Three Pareto Frontier results on Pythia-70M
(a) Sparsity vs. Unexplained
(b) Sparsity vs. L2 Loss
(c) Sparsity vs. Cosine Similarity
Figure 6: Three Pareto Frontier results on Pythia-160M
(a) Sparsity vs. Unexplained
(b) Sparsity vs. L2 Loss
(c) Sparsity vs. Cosine Similarity
Figure 7: Three Pareto Frontier results on Gemma-2-2B
5 Related Work
5.1 Sparse Autoencoders and Improvements
@@
A significant challenge in neural network interpretability is polysemanticity, where units (e.g., neurons) respond to diverse, semantically distinct inputs, complicating functional analysis (Elhage et al., 2022; Olah et al., 2020). The superposition hypothesis (Park et al., 2023) suggests that networks represent more features than available neurons by encoding them as directions in activation space, rather than solely via individual neuron activities. SAEs offer an unsupervised dictionary learning approach to tackle this, decomposing internal network activations (particularly in LLMs) to reveal latent interpretable units (Ferrando et al., 2024; Shu et al., 2025). Core to SAEs is an encoder mapping an activation $x$ to a higher-dimensional sparse representation $z$, and a decoder reconstructing $\hat{x}$ from $z$. Goal for SAEs are to isolate monosemantic and composable features, thus offering a more faithful representation of the model’s internal computational state (Cunningham et al., 2023; Huben et al., 2023).
@@
5.2 Linear Probe in Large Language Models
Linear probes have emerged as a fundamental method for elucidating how LLMs represent complex information within their activation spaces (Li et al., 2023; Von Rütte et al., 2024; Mikolov et al., 2013). This approach is grounded in the hypothesis that important high-level concepts are encoded linearly as directions in representational space (Kim et al., 2025; Liu et al., 2024). A substantial body of research supports the finding that linear probes are more effective than nonlinear probes at aligning model representations with specific behaviors. For instance, Gurnee and Tegmark (2023) applied linear probes to Llama-2 models to reveal that LLMs learn linear representations of space and time that are robust to prompting variations, unified across entity types, and encoded by specific neurons within the network. Similarly, Tigges et al. (2023) demonstrated that sentiment in language models emerges along specific linear directions in activation space, with a single dimension causally controlling sentiment polarity (positive versus negative) in model outputs through direct interventions. Additionally, concepts such as topic direction (Turner et al., 2023), political ideology (Kim et al., 2025), game states (Nanda et al., 2023), truthfulness (Marks and Tegmark, 2023) and safety (Arditi et al., 2024) have also been identified as important features that are linearly encoded within the internal activation spaces of LLMs.
6 Conclusions and Future Work
In this paper, we introduce AdaptiveK SAE, demonstrating that dynamically adjusting sparsity based on input complexity significantly improves representation decomposition in LLMs. By establishing that text complexity is linearly encoded in LLM activations, we developed a framework that allocates computational resources proportionally to content complexity, eliminating the need for extensive hyperparameter tuning while consistently outperforming fixed-sparsity baselines across multiple model scales. The experiments on Pythia-70M, Pythia-160M, and Gemma-2-2B confirm that this complexity-driven adaptation achieves better reconstruction fidelity, explained variance, and cosine similarity metrics. In future, we plan to explore applying AdaptiveK SAE to multimodal models, and investigate how our adaptive feature allocation might enhance interpretability for safety alignment.
Appendix A Linear Probes Training Details
@@
Sample 1:
True Complexity: 1.2 $|$ Pred Complexity: 3.18 $|$ K value: 95 $|$ Activated Features: 95
@@
Sample 2:
True Complexity: 3.5 $|$ Pred Complexity: 4.26 $|$ K value: 137 $|$ Activated Features: 137
@@
Sample 3:
True Complexity: 5.2 $|$ Pred Complexity: 5.39 $|$ K value: 187 $|$ Activated Features: 187
@@
Sample 4:
True Complexity: 6.2 $|$ Pred Complexity: 6.05 $|$ K value: 216 $|$ Activated Features: 216
@@
Sample 5:
True Complexity: 9.1 $|$ Pred Complexity: 9.19 $|$ K value: 298 $|$ Activated Features: 242
parameters of the MSSM and to trace back, sector-wise, the sensitivity to initial conditions of the Yukawa couplings and the soft susy breaking parameters. We have established analytically a generic screening of non-universality, in the vicinity of the infrared quasi fixed points. In practice, this property gives the general trend of the behaviour, despite the large number of free parameters, and even when one is not very close to such a quasi fixed point. This shows that non-universality of th...
Sample 6:
True Complexity: 9.6 $|$ Pred Complexity: 8.91 $|$ K value: 294 $|$ Activated Features: 294
@@
Appendix B Additional SAE Training Details
Our SAEs are trained on the residual stream because researchers typically focus on this component when interpreting or steering model behaviors. This alignment ensures the learned representations directly support the most common SAE applications in model analysis and intervention. Our code is available at: https://anonymous.4open.science/r/adaptiveK-5258
@@ Appendix B Additional SAE Training Details
SAE Sparsity Pythia Gemma
@@
Matryoshka Batch TopK 20, 40, 80, 160, 320, 640
Gated Sparsity Penalities 0.6, 0.9, 1.2, 2, 3, 4 2.4, 3.6, 4.8, 8, 12, 16
Relu 0.6, 0.9, 1.2, 2, 3, 4 2.4, 3.6, 4.8, 8, 12, 16
Relu_new 0.6, 0.9, 1.2, 2, 3, 4 2.4, 3.6, 4.8, 8, 12, 16
P Anneal 0.3, 0.45, 0.6, 1, 1.5, 2 1.2, 1.8, 2.4, 4, 6, 8
Appendix C Additional Experimental Evaluation
C.1 Analysis of Layers
@@
We trained AdaptiveK SAE on each layer of Pythia-160M and on layers 4, 8, 12, 16, 20, and 24 of the Gemma-2-2B model. A cross-layer comparison of key performance metrics is presented in Table 3 and 4. L2 ratio measures the proportion between the L2 norms of reconstructed and original activations, with values closer to 1 indicating preservation of the original activation magnitude. AdaptiveK exhibits robust performance across all tested layers in both models, with Explained Variance, Cosine Similarity, and L2 Ratio consistently above 0.74, 0.91, and 0.89 respectively. This confirms the algorithm’s generalizability throughout the entire LLM hierarchy.
Table 3: Layer-wise performance in Pythia-160M
@@
Layer Explained Variance Cosine Similarity L2 Ratio
0 0.90 0.96 0.91
1 0.86 0.95 0.91
2 0.85 0.95 0.91
3 0.84 0.94 0.89
4 0.83 0.94 0.89
5 0.85 0.96 0.94
6 0.89 0.96 0.93
7 0.91 0.97 0.95
8 0.91 0.97 0.96
9 0.89 0.96 0.95
10 0.86 0.96 0.95
11 0.89 0.99 0.99
Table 4: Layer-wise performance in Gemma-2-2B
Layer Explained Variance Cosine Similarity L2 Ratio
4 0.81 0.93 0.93
8 0.80 0.93 0.93
12 0.74 0.91 0.92
16 0.77 0.92 0.93
20 0.79 0.93 0.94
24 0.77 0.92 0.93
C.2 Other Evaluation Metrics

### v2

@@ 4.3 Pareto Frontier Results
We evaluated three different metrics vs. sparsity frontier for benchmark SAEs with different sparsity constraints. All SAEs have a dictionary size of 16384. These three metrics are:
@@
  • Reconstruction L2 Loss: $\|x-\hat{x}\|_{2}^{2}$ averages the squared Euclidean distance between each reconstruction and its input, lower is better.
@@
  • Fraction of Variance Explained: $\frac{\mathrm{Var}(x-\hat{x})}{\mathrm{Var}(x)}$ measures how much input variability is captured by the reconstruction. The plots show 1 minus this value (unexplained variance, lower is better).
@@
  • Cosine Similarity: $\frac{x\cdot\hat{x}}{\|x\|_{2}\|\hat{x}\|_{2}}$ evaluates directional fidelity between original and reconstructed vectors. The plots show 1 minus this value, thus lower is better.
[…]
@@
Notably, across all model scales, AdaptiveK’s reconstructed activations more accurately match the originals in both distance and direction (Figure 5, 12, 7, 14) while explaining a higher proportion of data variance (Figure 6, 13), demonstrating AdaptiveK’s robustness and generalizability.
4.4 Interpretability Analysis of AdaptiveK SAE
4.4.1 Definition of LLM Interpretability
Following established mechanistic interpretability research (Rai et al., 2025; Shu et al., 2025; Zhao et al., 2023), we define LLM interpretability as how well learned features correspond to semantically coherent, human-understandable concepts that can be consistently identified across different contexts. Specifically, we consider a feature to be interpretable if it satisfies: (1) input-based semantic coherence: it activates on inputs sharing a clear conceptual relationship, and (2) output-based human comprehensibility: it influences semantically consistent vocabulary prediction.
4.4.2 Measures of Interpretability
For input-based analysis, we implemented the MaxAct (Bricken et al., 2023) method to identify text segments that maximally activate each SAE feature. Specifically, for each target feature $w_{m}$, we fed corpus texts $x$ from our test set into the original language model to obtain hidden representations $\mathbf{h}_{x}=f_{<l}(x)$ at layer $l$, computed feature activation strengths $a_{m}(x)=\operatorname{ReLU}\left(\mathbf{w}_{m}^{\top}\cdot\mathbf{h}_{x}+b_{m}\right)$ through the SAE encoder, and selected the top-$K$ text segments with highest activations:
$\mathcal{I}_{m}=\underset{\left|\mathcal{X}^{\prime}\right|=K}{\arg\max}\sum_{x\in\mathcal{X}^{\prime}}a_{m}(x)$ (11)
Through experiments on Gemma-2-9B comparing against TopK SAE (k=80), we analyzed thematic consistency of high-activation texts to quantify whether features learned monosemantic, human-interpretable concepts. Figure 8 shows that for biomedical texts, AdaptiveK’s activation patterns focus precisely on core professional concepts, including technical terms (“resonance spectroscopy,” “MRS,” “ppm”), biochemical concepts (“acetylcarnitine,” “peak”), and methodological vocabulary (“verification,” “chemical,” “induction”). In contrast, TopK activates both similar professional terminology and numerous semantically weak function words (“been,” “groups,” “generated,” “aimed”), diluting feature focus and inefficient computational resource utilization with diminishing semantic returns despite employing more features. This demonstrates AdaptiveK’s ability to identify semantic complexity and allocate appropriate high-quality features while avoiding irrelevant activations.
Figure 8: Input-based interpretability analysis using MaxAct method.
For output-based analysis, we used VocabProj (Shu et al., 2025) to analyze feature influence patterns on model vocabulary prediction. By computing inner products between each feature vector $w_{m}$ in the SAE decoder and the language model’s output vocabulary embedding matrix $f_{out}(w)$, we obtained logits quantify which words model tends to generate or suppress when features activate:
$\mathcal{I}_{m}=\arg\max_{w\in\mathcal{V}}f_{\text{out }}(w)\cdot\mathbf{w}_{m}^{\top}$ (12)
We extracted the top 10 positive and negative associated vocabulary items per feature to evaluate semantic coherence across Llama-3.1-8B and Gemma-2-9B (Figure 9). In Gemma-2-9B, Feature 5637 specializes in “title” concepts, with positive vocabulary centered around title variants including different capitalizations and languages (e.g., German “Titel”), while negative vocabulary contains unrelated noun. Feature 2370 focuses on “abstract” concepts, similarly demonstrating cross-linguistic and cross-format unity with variants like Spanish “abstracto”, and negative vocabulary containing technical symbols and irrelevant words. In Llama-3.1-8B, Feature 2865 concentrates on “microscopic” concepts, with positive vocabulary extending from “micro” to related terms like “microscopic” and “minuscule,” showing semantic expansion from root to concept. Feature 8949 captures “expansion” across tenses and noun forms, reflecting complete semantic field coverage.
Figure 9: Output-based interpretability analysis using VocabProj method.
These results reveal AdaptiveK’s capability for learning semantically coherent features capturing not only literal expressions but also concept variants across contexts, languages, and grammatical forms. The negative vocabulary ranging from foreign terms to technical terminology consistently forms clear semantic contrasts with target concepts, indicating robust conceptual boundary identification.
5 Conclusion
In this paper, we introduce AdaptiveK SAE, which demonstrates that dynamically adjusting sparsity based on input complexity significantly improves representation decomposition in LLMs. By establishing that text complexity is linearly encoded in LLM activations, we developed a framework that allocates computational resources proportionally to content complexity, eliminating extensive hyperparameter tuning while consistently outperforming fixed-sparsity baselines across multiple model scales. Experiments on 8 LLMs confirm that complexity-driven adaptation achieves better reconstruction fidelity, explained variance, cosine similarity and interpretability.
Reproducibility Statement
We provide comprehensive resources to ensure the reproducibility of our work. Our code implementation, including the complete AdaptiveK SAE framework, linear probe training pipeline, and evaluation scripts, is available at https://anonymous.4open.science/r/adaptiveK-5258. The training dataset consists of 250,000 contexts from pile-uncopyrighted (Gao et al., 2020), with complexity annotations generated using GPT-4.1-mini based on six semantic dimensions detailed in Appendix B. All training details and hyperparameters are specified in Section 3.3 and Appendix F.4. We trained on the last token representations of 1024-token contexts using Adam optimizer with learning rate 1e-3. The three-phase training procedure and adaptive deviation penalty mechanism are fully described in Algorithm 1. All baseline SAE implementations were obtained from the dictionary_learning library (Marks et al., 2024b), with their specific configurations listed in Table 2.
Appendix A Related Work
A.1 Sparse Autoencoders and Improvements
@@
A significant challenge in neural network interpretability is polysemanticity, where units (e.g., neurons) respond to diverse, semantically distinct inputs, complicating functional analysis (Elhage et al., 2022; Olah et al., 2020). The superposition hypothesis (Park et al., 2023) suggests that networks represent more features than available neurons by encoding them as directions in activation space, rather than solely via individual neuron activities. SAEs offer an unsupervised dictionary learning approach to tackle this, decomposing internal network activations (particularly in LLMs) to reveal latent interpretable units (Ferrando et al., 2024; Shu et al., 2025). Core to SAEs is an encoder mapping an activation $x$ to a higher-dimensional sparse representation $z$, and a decoder reconstructing $\hat{x}$ from $z$. Goal for SAEs are to isolate monosemantic and composable features, thus offering a more faithful representation of the model’s internal computational state (Huben et al., 2023).
@@
A.2 Linear Probe in Large Language Models
Linear probes have emerged as a fundamental method for elucidating how LLMs represent complex information within their activation spaces (Li et al., 2023; Von Rütte et al., 2024; Mikolov et al., 2013). This approach is grounded in the hypothesis that important high-level concepts are encoded linearly as directions in representational space (Kim et al., 2025; Liu et al., 2024). A substantial body of research supports the finding that linear probes are more effective than nonlinear probes at aligning model representations with specific behaviors. For instance, Gurnee & Tegmark (2023) applied linear probes to Llama-2 models to reveal that LLMs learn linear representations of space and time that are robust to prompting variations, unified across entity types, and encoded by specific neurons within the network. Similarly, Tigges et al. (2023) demonstrated that sentiment in language models emerges along specific linear directions in activation space, with a single dimension causally controlling sentiment polarity (positive versus negative) in model outputs through direct interventions. Additionally, concepts such as topic direction (Turner et al., 2023), political ideology (Kim et al., 2025), game states (Nanda et al., 2023), truthfulness (Marks & Tegmark, 2023) and safety (Arditi et al., 2024) have also been identified as important features that are linearly encoded within the internal activation spaces of LLMs.
Appendix B Linear Probes Training Details
@@
Sample 5: True Complexity: 9.1, Pred Complexity: 9.19, K value: 298, Activated Features: 242 parameters of the MSSM and to trace back, sector-wise, the sensitivity to initial conditions of the Yukawa couplings and the soft susy breaking parameters. We have established analytically a generic screening of non-universality, in the vicinity of the infrared quasi fixed points. In practice, this property gives the general trend of the behaviour, despite the large number of free parameters, and even when one is not very close to such a quasi fixed point. This shows that non-universality of th...
@@
Appendix C Additional SAE Training Details
Our SAEs are trained on the residual stream because researchers typically focus on this component when interpreting or steering model behaviors. This alignment ensures the learned representations directly support the most common SAE applications in model analysis and intervention.
@@ Appendix C Additional SAE Training Details
SAE Sparsity Pythia Gemma/Llama/Qwen/Phi
@@
Matryoshka 20, 40, 80, 160, 320, 640
Gated Sparsity Penalities 0.6, 0.9, 1.2, 2, 3, 4 0.012, 0.015, 0.02, 0.03, 0.04, 0.06
Relu 0.6, 0.9, 1.2, 2, 3, 4 0.012, 0.015, 0.02, 0.03, 0.04, 0.06
Relu_new 0.6, 0.9, 1.2, 2, 3, 4 0.012, 0.015, 0.02, 0.03, 0.04, 0.06
P Anneal 0.3, 0.45, 0.6, 1, 1.5, 2 0.006, 0.008, 0.01, 0.015, 0.02, 0.025
Appendix D More Results
D.1 Linear Probe Performance
Figure 10 presents linear probe performance across a broader set of LLMs.
(a) Pythia-160M
(b) Gemma-2-9B
(c) Llama-3.1-8B
(d) Qwen-3-14B
Figure 10: Supplement to the results of Figure 3
D.2 Relationship between Complexity Scores and Allocated Feature Counts
Figure 11 presents the relationship between complexity scores and allocated feature counts across a broader set of LLMs.
(a) Pythia-160M
(b) Gemma-2-9B
(c) Llama-3.1-8B
(d) Qwen-3-14B
Figure 11: Supplement to the results of Figure 4
D.3 Pareto Frontier Results
Figure 12, 13 and 14 presents L2 Loss, Unexplained Variance and Cosine Similarity pareto frontier results across a broader set of LLMs.
(a) Pythia-160M
(b) Gemma-2-9B
(c) Llama-3.1-8B
(d) Qwen-3-14B
Figure 12: Supplement to the results of Figure 5
(a) Pythia-160M
(b) Gemma-2-9B
(c) Llama-3.1-8B
(d) Qwen-3-14B
Figure 13: Supplement to the results of Figure 6
(a) Pythia-160M
(b) Gemma-2-9B
(c) Llama-3.1-8B
(d) Qwen-3-14B
Figure 14: Supplement to the results of Figure 7
Appendix E Additional Linear Probe Evaluation
E.1 PCA Dimensionality Reduction Experiments
To provide more rigorous statistical validation, PCA dimensionality reduction experiments were conducted on the activation data. Specifically, since the original probe dimensionality is 2048 (for Gemma-2-2B), it could theoretically ”memorize” substantial information. However, if complexity is truly linearly encoded, then only a few dimensions should be needed for accurate prediction.
We performed PCA decomposition on all context activation vectors, identifying the top k directions with maximum variance (principal components), and projected the original 2048-dimensional activations onto these k directions (k=10-500). Probes for complexity prediction were then trained on the low-dimensional representations. Figure 15 shows the RMSE and Pearson correlation coefficients of probes using different numbers of principal components. Using just 200 principal components (9.8% of full dimensionality), the probe achieves a Pearson correlation of 0.72 compared to 0.79 for the full-dimensional probe, recovering 91% of the predictive performance while explaining 75.6% of the variance. With 400 components (19.5% of dimensions), performance reaches 96% of the full probe (Pearson: 0.76 vs 0.79) with 88.4% explained variance. This demonstrates that text complexity is indeed linearly encoded. If it were merely high-dimensional memorization, many more principal components would be required to achieve good predictive performance.
(a) RSME vs. PCA dimensions
(b) Pearson vs. PCA dimensions
(c) Exp Var vs. PCA dimensions
Figure 15: PCA dimensionality reduction experiments on Gemma-2-2B
E.2 Layer-wise Evaluation
To further validate the effectiveness of complexity prediction using linear probes, we conducted layer-wise experiments across all 26 layers of Gemma-2-2B to understand how complexity representations develop throughout the model. Following previous work (Gurnee & Tegmark, 2023), which found that probing performance exhibits a characteristic pattern of initial growth followed by saturation as layer depth increases, our experimental results in Figure 16 demonstrate the same trajectory. Starting from layer 4 with a Pearson correlation of 0.766, performance steadily improves through the middle layers, reaching peak performance at layer 22 with a correlation of 0.814 and RMSE of 1.18. Beyond this point, performance plateaus and even slightly decreases (layer 24: 0.801). This demonstrates that LLM representations contain complexity information in the same way they contain spatial and temporal information, and that deeper layers are progressively better at capturing text complexity.
Figure 16: Layer-wise evaluation of linear probe on Gemma-2-2B
Appendix F Additional SAE Evaluation
F.1 Layer-wise Evaluation
@@
We trained AdaptiveK SAE on each layer of Pythia-160M and on layers 4, 8, 12, 16, 20, and 24 of the Gemma-2-2B model. A cross-layer comparison of key performance metrics is presented in Figure 17 and 18. L2 ratio measures the proportion between the L2 norms of reconstructed and original activations, with values closer to 1 indicating preservation of the original activation magnitude. AdaptiveK exhibits robust performance across all tested layers in both models, with Explained Variance, Cosine Similarity, and L2 Ratio consistently above 0.74, 0.91, and 0.89 respectively. This confirms the algorithm’s generalizability throughout the entire LLM hierarchy.
(a) Layer-wise L2 Ratio
(b) Layer-wise Cosine Similarity
(c) Layer-wise Explained Variance
Figure 17: Layer-wise performance on Pythia-160M
(a) Layer-wise L2 Ratio
(b) Layer-wise Cosine Similarity
(c) Layer-wise Explained Variance
Figure 18: Layer-wise performance on Gemma-2-2B
F.2 Extending to Encoder-only and Encoder-decoder Models
We extended our model to encoder-only BERT and encoder-decoder T5, with results shown in Table 3. BERT-340M achieved 0.89 cosine similarity and 0.82 L2 ratio, demonstrating successful adaptation to bidirectional attention mechanisms. T5-small results show the encoder outperforms the decoder by 23.56% in explained variance, indicating that input understanding tasks are relatively regular while generation tasks are more complex. The encoder also performs better on other metrics, suggesting that encoders focus on understanding input semantic representations while decoders must handle both understanding and generation tasks simultaneously. All results maintain good reconstruction quality (cosine similarity more than 0.89, L2 ratio close to 1.0), proving the cross-architecture universality of “complexity → more features”.
Table 3: Performance of encoder-only and encoder-decoder models
@@ F.2 Extending to Encoder-only and Encoder-decoder Models
Model Layer SAE Position Explained Variance Cosine Similarity L2 Ratio
BERT-340M 8 encoder 0.66 0.89 0.8225
T5-small 3 encoder 0.97 0.98 1.0093
T5-small 3 decoder 0.74 0.97 1.0139
F.3 Other Evaluation Metrics